OpenStudy (anonymous):

The sum of a negative number and its square is 110. Find the number.

OpenStudy (isaiah.feynman):

Let the number be n.

OpenStudy (isaiah.feynman):

Solve for n.

OpenStudy (anonymous):

im so lost

OpenStudy (isaiah.feynman):

Where are you lost?

OpenStudy (princeharryyy):

-n + (-n)^2 =110 n^2-n-110 = 0 (n-11)(n+10)=0; => n=11 or n = -10

OpenStudy (anonymous):

thank you @princeharryyy i have no idea how to do this

OpenStudy (princeharryyy):

do u have it now @hislightningbug812 this was easy...

OpenStudy (anonymous):

yea i got it thanx

OpenStudy (princeharryyy):

u should get a medal too. here it's for u... happy learning.. )))

OpenStudy (princeharryyy):

@hislightningbug812

OpenStudy (anonymous):

thanx :)

OpenStudy (princeharryyy):

u r wlcm @hislightningbug812

OpenStudy (triciaal):